The variable x is 4 less than the variable y.
1. If the variable t is multiplied by 5, the result is equal to the variable R.
2. If 3 is subtracted from the product of the variables x and y, the result is equal to 4 times the variable z.
3. The quotient of the variable x with the variable y is 2 more than 0.7 of the variable z.

Answer by stanbon(48545)

The variable x is 4 less than the variable y.
x=y-4
-------------
1. If the variable t is multiplied by 5, the result is equal to the variable R.
5t=R
----------
2. If 3 is subtracted from the product of the variables x and y, the result is equal to 4 times the variable z.
(xy)-3 = 4z
-------------
3. The quotient of the variable x with the variable y is 2 more than 0.7 of the variable z.
(x/y)=0.7z+2
======================
Cheers,
Stan H.